OpenStudy (anonymous):

i had did this and i didnt get a good grade on this and i dont understand why can somebody help me with this please Launch Area:___(1, 2)___ Point A:____(4,5)______ Point B:__(4,3)________ Point C:__(5,-5)________ You must show your work on each question below. 1. Determine the equation of the line, in standard form, that will get your spacecraft from the Launch Area to Point A. 2. Determine the equation of the line, in point-slope form, that will get your spacecraft from Point A to Point B.

OpenStudy (anonymous):

Post what you did so I can see where you went wrong.

OpenStudy (anonymous):

1. Determine the equation of the line, in standard form, that will get your spacecraft from the Launch Area to Point A. -Y2-Y1/X2-X1 5-2/4-1=3/3=1 y-y1=m(x-x1) y-2=11(x+1) y-2+1x+1 y=1x+3 -1x+y=3 2. Determine the equation of the line, in point-slope form, that will get your spacecraft from Point A to Point B. - Y2-Y1/X2-X1 3-5/4-4= -2/0 y-y1=m(x-x1) y-4=-2(x-5) y-4=-2x+10 2x+y-4=10 2x+y=14

OpenStudy (anonymous):

and here are the other two problems i got wrong too 3. Determine the equation of the line, in slope-intercept form, that will get your spacecraft from Point B to Point C. 4. In question 2, you selected one of two points (Point A or Point B) to be included in your point-slope equation. Write the point-slope form of that equation again, using the other point’s coordinates. 5. Convert the equations you arrived at in question 2 and question 4 into slope-intercept form.

OpenStudy (anonymous):

If I am correct... launch area = Starting point point A = Ending destination  #2 slope recheck numbers